# Modulo arithmetic of function spaces: Subset hyperspaces as quotients of function spaces

### v1

@@ 3. Subset hyperspaces as quotients of function spaces
    $\tau_{v}=\langle\tau_{pq}\cup\tau_{gq}\rangle\subset\langle\tau_{p}\cup\tau_{s}\rangle_{q}$,  where if $X$ is $T_{1}$, then equality holds iff $Y$ is finite (in which case $\tau_{v}=\tau_{pq}$).
@@
By Lemmas 3.14 and 3.16, we get both (1) via $\tau_{v}=\langle\tau_{v}^{-}\cup\tau_{v}^{+}\rangle$ and (2) via $q^{-1}(\tau_{v})=q^{-1}(\langle\tau_{v}^{-}\cup\tau_{v}^{+}\rangle)=\langle q^{-1}(\tau_{v}^{-})\cup q^{-1}(\tau_{v}^{+})\rangle$. ∎
Remark 3.18.
@@
    In the above results for $q:(\mathcal{F},\tau)\rightarrow(Cl_{Y}(X),\tau_{q})$ with $\tau_{q}=\tau_{v}$, if $FS_{n}(X)\subset(Cl_{Y}(X),\tau_{q})$ is closed, then (by Remark 2.3(3)) $q:(q^{-1}(FS_{n}(X))\cap\mathcal{F},\tau)\subset(\mathcal{F},\tau)\rightarrow(FS_{n}(X),\tau_{q})$ is a quotient map as well. In this case, Theorem 3.17(1) holds (i) with $\mathcal{F}$ and $Cl_{Y}(X)$ replaced by $q^{-1}(FS_{n}(X))\cap\mathcal{F}$ and $FS_{n}(X)$ respectively, and (ii) with equality for any $Y$.
@@
    If $(\mathcal{F},\tau)$ is compact, connected, or path-connected then so is $(Cl_{Y}(X),\tau_{q})$, since $q:(\mathcal{F},\tau)\rightarrow(Cl_{Y}(X),\tau_{q})$ is continuous.
@@
    Continuous maps $h:(Cl_{Y}(X),\tau_{q})\rightarrow Z$ (for a space $Z$) are precisely continuous maps $h:(\mathcal{F},\tau)\rightarrow Z$ that are constant on the equivalence classes $[f]:=q^{-1}(q(f))=\left\{g\in\mathcal{F}:q(g)=q(f)\right\}$ (i.e., $h|_{[f]}=const$) for all $f\in\mathcal{F}$.
@@
    Let $Y$ be a manifold, $X=\bigsqcup_{y\in Y}X_{y}=\bigsqcup_{y\in Y}\pi^{-1}(y)$ a fiber bundle over $Y$ with projection $\pi:X\rightarrow Y$. Let $\mathcal{X}:=\{X_{y}\}_{y\in Y}=\{\pi^{-1}(y)\}_{y\in Y}$. Then, in the notation of Definition 3.1, we get a map on global sections of $Y$ given by
@@
    In this case, the limit topology of Definition 3.1 on $X$ plays a nontrivial role as a topology (induced by that of the fibers $X_{y}$ of the bundle $X$) that can be compared with the underlying topology of $X$.
@@ 4. Concrete quotient-realization and metrization of compact-subset hyperspaces
Definition 4.1 (Chained space, Topologies of uniform/pointwise/compact convergence).
A chained-space $X=(X,\mathcal{U})$ is a space $X$ together with a chain $\mathcal{U}=(\{\mathcal{U}_{\lambda}\}_{\lambda\in\Lambda},\preceq)$ of open covers of $X$, where “$\mathcal{U}_{\lambda}\preceq\mathcal{U}_{\lambda^{\prime}}$ iff for any $O\in\mathcal{U}_{\lambda}$ there is $O^{\prime}\in\mathcal{U}_{\lambda^{\prime}}$ such that $O\subset O^{\prime}$ and [for any $O^{\prime}\in\mathcal{U}_{\lambda^{\prime}}$ and any $O\in\mathcal{U}_{\lambda}$, $O^{\prime}$ is not a proper subset of $O$]”. A collection of open sets $\mathcal{O}$ of $X$ is homogeneous (with respect to $\mathcal{U}$) if $\mathcal{O}\subset\mathcal{U}_{\lambda}$ for some $\lambda$.
[…]
NB: All subsequent concepts/results based on (conditional-) uniform convergence, that do not refer (hence do not depend on) any details about the chain $\mathcal{U}$, also apply-to/hold-for unconditional-uniform convergence. In particular, Notes 1-3 below (and more) do not depend on $\mathcal{U}$.
@@
Note 1: If $Z\subset Z^{\prime}\subset Y$, then $f_{\alpha}\stackrel{{\scriptstyle u|_{Z^{\prime}}}}{{\longrightarrow}}f$ implies $f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f$ (which holds because each system of open sets $\mathcal{O}:=\{O_{z}\ni f(z):O_{z}\in\mathcal{U}_{\lambda},z\in Z\}$ can be extended to a system of open sets $\mathcal{O}^{\prime}:=\{O_{z}\ni f(z):O_{z}\in\mathcal{U}_{\lambda},z\in Z^{\prime}\}$). In particular, for any $Z\subset Y$,
[…]
@@
A net $\{f_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{F}$ converges uniformly wrt a family of sets $\mathcal{Z}\subset\mathcal{P}^{\ast}(Y)$ to $f\in\mathcal{F}$ (making it a $\mathcal{Z}$-uniformly convergent net), written $f_{\alpha}\stackrel{{\scriptstyle u|_{\mathcal{Z}}}}{{\longrightarrow}}f$, if $f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f$ for all $Z\in\mathcal{Z}$. We note that $\{f_{\alpha}\}$ is uniformly convergent if and only if $\{Y\}$-uniformly convergent, if and only if (by Note 1 above) uniformly convergent on $Z$ for every $Z\subset Y$.
@@
A subset $\mathcal{C}\subset X^{Y}$ is $\mathcal{Z}$-uniformly closed if every $\mathcal{Z}$-uniformly convergent net in $\mathcal{C}$ converges to a point in $\mathcal{C}$. The topology of $\mathcal{Z}$-uniform convergence $\tau_{uc|_{\mathcal{Z}}}$ on $X^{Y}$ is the topology whose closed sets are the $\mathcal{Z}$-uniformly closed subsets of $X^{Y}$. When $\mathcal{Z}=\{Y\}$, we simply drop $\mathcal{Z}$ from the terminology, i.e., “$\mathcal{Z}$-uniformly closed”, $\tau_{uc|_{\mathcal{Z}}}$ (“topology of $\mathcal{Z}$-uniform convergence”), etc become “uniformly closed”, $\tau_{uc}$ (“topology of uniform convergence”), etc. If $\mathcal{Z}=\big{\{}\{y\}:y\in Y\big{\}}$, then $\tau_{uc|_{\mathcal{Z}}}=\tau_{p}$, in which case we replace “uniform” or “uniformly” with “pointwise”. When $Y$ is a space and $\mathcal{Z}=K(Y)$, we call $\tau_{\textrm{cc}}:=\tau_{uc|_{K(Y)}}$ the topology of compact convergence.
@@
Note 2: If $\mathcal{Z}\subset\mathcal{Z}^{\prime}$, then $\tau_{uc|_{\mathcal{Z}}}\subset\tau_{uc|_{\mathcal{Z}^{\prime}}}$. To see this, take a $\tau_{uc|_{\mathcal{Z}}}$-closed set $\mathcal{C}\subset X^{Y}$ and show that it is also $\tau_{uc|_{\mathcal{Z}^{\prime}}}$-closed: Indeed, if $f_{\alpha}\in\mathcal{C}$, then $f_{\alpha}\stackrel{{\scriptstyle\tau_{uc|_{\mathcal{Z}^{\prime}}}}}{{\longrightarrow}}f$ implies $f_{\alpha}\stackrel{{\scriptstyle\tau_{uc|_{\mathcal{Z}}}}}{{\longrightarrow}}f$, and so $f\in\mathcal{C}$.).
@@
Note 3: For any $\mathcal{Z}\subset\mathcal{P}(Y)$, we have ${}\tau_{uc|_{\mathcal{Z}}}=\tau_{uc|_{\textrm{UP}(\mathcal{Z})}},~{}$ where ${}\textrm{UP}(\mathcal{Z}):=\bigcup_{Z\in\mathcal{Z}}\mathcal{P}(Z)$. This follows from Note 1 above.
@@
Note 4: If $X$ is a metric space, $Y$ a space, and $C(Y,X)\subset X^{Y}$ continuous maps, then $f_{\alpha}\stackrel{{\scriptstyle\tau_{cc}}}{{\longrightarrow}}f$ in $C(Y,X)$ if and only if $f_{\alpha}\stackrel{{\scriptstyle d^{K}}}{{\longrightarrow}}f$ (i.e., $\forall\varepsilon>0$, $\exists\alpha^{\varepsilon}$ s.t. $d^{K}(f_{\alpha},f)<\varepsilon$ $\forall\alpha\geq\alpha^{\varepsilon}$) in $C(Y,X)$ for every compact set $K\in K(Y)$, (footnote999 Proof: ($\Rightarrow$): Assume $f_{\alpha}\stackrel{{\scriptstyle\tau_{cc}}}{{\longrightarrow}}f$. Fix any $K\in K(Y)$. Then for any collection $\mathcal{O}=\{O\ni f(y):y\in K\}$ of open sets in $X$, there is $\alpha^{\mathcal{O}}$ such that $\{f_{\alpha}(y)\}_{\alpha\geq\alpha^{\mathcal{O}}}\subset O_{y}$ (for all $y\in K$). So, for any $\varepsilon>0$, if $O_{y,\varepsilon}:=B_{\varepsilon}^{K}(f(y))$ and $\mathcal{O}_{\varepsilon}:=\{O_{y,\varepsilon}:y\in K\}$, then $d^{K}(f_{\alpha},f)=\sup_{y\in K}d(f_{\alpha}(y),f(y))<2\varepsilon~{}~{}~{}~{}\textrm{for all}~{}~{}\alpha\geq\alpha^{\varepsilon}:=\alpha^{\mathcal{O}_{\varepsilon}}$. ($\Leftarrow$): Assume $f_{\alpha}\stackrel{{\scriptstyle d^{K}}}{{\longrightarrow}}f$ in $C(Y,X)$ for all $K\in K(Y)$. Fix $K\in K(Y)$. For any $\varepsilon>0$, there is $\alpha^{\varepsilon}$ such that $d^{K}(f_{\alpha},f)<\varepsilon$ for all $\alpha\geq\alpha^{\varepsilon}$. Therefore, $\{f_{\alpha}(y)\}_{\alpha\geq\alpha^{\varepsilon}}\subset B_{\varepsilon}(f(y))$ for all $\in K$. Consider any collection $\mathcal{O}=\{O_{y}\ni f(y):y\in K\}$ of open sets in $X$. Let $B_{r_{y}}(f(y))\subset O_{y}$ for each $y\in K$. Then by the compactness of $f(K)$, we can choose $r>0$ (independent of $y$) such that $B_{r}(f(y))\subset O_{y}$ for each $y\in K$. Hence, with $\varepsilon<r$, we get $\{f_{\alpha}(y)\}_{\alpha\geq\alpha^{\varepsilon}}\subset B_{\varepsilon}(f(y))\subset B_{r}(f(y))\subset O_{y}$ for all $\in K$. ), where $d^{K}(f_{\alpha},f):=\sup_{y\in K}d(f_{\alpha}(y),f(y))$.
@@
$\mathcal{S}\mathcal{B}_{luc}:=\left\{B^{d^{K}}_{r}(f):r>0,f\in\mathcal{F},K\in K(Y)\right\}$
(which, by Note 4 above, coincides with the topology of compact convergence $\tau_{cc}$ on continuous functions $\mathcal{F}:=C(Y,X)$).
Note 5: On any $\mathcal{F}\subset X^{Y}$,   $\tau_{cc}\subset\tau_{luc}$.
Definition 4.2 (An interpretation of the Vietoris topology).
Let $X$ be a space and $A_{\alpha},A\in\mathcal{P}^{\ast}(X)$. We say $A_{\alpha}$ fully-converges (resp., partially-converges) to $A$ in $X$, written $A_{\alpha}\stackrel{{\scriptstyle+}}{{\longrightarrow}}A$ (resp., $A_{\alpha}\stackrel{{\scriptstyle-}}{{\longrightarrow}}A$), if for any open set $O\subset X$ containing $A$ (resp., meeting $A$), there exists $\alpha_{+}^{O}$ (resp., $\alpha_{-}^{O}$) such that
@@
that is, if $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}^{+}}}{{\longrightarrow}}A$ (resp., $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}^{-}}}{{\longrightarrow}}A$). A set $\mathcal{C}\subset P^{\ast}(X)$ is fully-closed (resp., partially-closed) if for any net $A_{\alpha}\in\mathcal{C}$ such that $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}^{+}}}{{\longrightarrow}}A$ (resp., $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}^{-}}}{{\longrightarrow}}A$), we have $A\in\mathcal{C}$. The upper Vietoris topology $\tau_{v}^{+}$ (resp., lower Vietoris topology $\tau_{v}^{-}$) of $\mathcal{P}^{\ast}(X)$ is the topology whose closed sets are the fully-closed (resp., partially-closed) subsets of $\mathcal{P}^{\ast}(X)$.

### v2

@@ 3. Subset hyperspaces as quotients of function spaces
For a finite $Y$, the following theorem (Theorem 3.3) realizes $\tau_{v}$ as a quotient of a $\tau_{p}$-compatible topology. For a general $Y$, the realization of $\tau_{v}$ as a of a $\tau_{p}$-compatible topology will be accomplished in Theorem 5.1 (which requires Theorem 3.3).
@@
    $\tau_{v}=\langle\tau_{pq}\cup\tau_{sq}\rangle\subset\langle\tau_{p}\cup\tau_{s}\rangle_{q}$,  where if $X$ is $T_{1}$, then equality holds iff $Y$ is finite (in which case $\tau_{v}=\tau_{pq}$).
@@
By Lemmas 3.1 and 3.2, we get both (1) via $\tau_{v}=\langle\tau_{v}^{-}\cup\tau_{v}^{+}\rangle$ and (2) via $q^{-1}(\tau_{v})=q^{-1}(\langle\tau_{v}^{-}\cup\tau_{v}^{+}\rangle)=\langle q^{-1}(\tau_{v}^{-})\cup q^{-1}(\tau_{v}^{+})\rangle$. ∎
Remark.
@@
    In the above results for $q:(\mathcal{F},\tau)\rightarrow(Cl_{Y}(X),\tau_{q})$ with $\tau_{q}=\tau_{v}$, if $FS_{n}(X)\subset(Cl_{Y}(X),\tau_{q})$ is closed, then (by Definition 2’s Remark (3)) $q:(q^{-1}(FS_{n}(X))\cap\mathcal{F},\tau)\subset(\mathcal{F},\tau)\rightarrow(FS_{n}(X),\tau_{q})$ is a quotient map as well. In this case, Theorem 3.3(1) holds (i) with $\mathcal{F}$ and $Cl_{Y}(X)$ replaced by $q^{-1}(FS_{n}(X))\cap\mathcal{F}$ and $FS_{n}(X)$ respectively, and (ii) with equality for any $Y$.
@@
    If $(\mathcal{F},\tau)$ is compact, connected, or path-connected, then so is $(Cl_{Y}(X),\tau_{q})$, since $q:(\mathcal{F},\tau)\rightarrow(Cl_{Y}(X),\tau_{q})$ is continuous.
@@
    Continuous maps $h:(Cl_{Y}(X),\tau_{q})\rightarrow Z$ (for a space $Z$) are precisely continuous maps $h:(\mathcal{F},\tau)\rightarrow Z$ that are constant on the equivalence classes $f]:=q^{-1}(q(f))=\left\{g\in\mathcal{F}:q(g)=q(f)\right\}$ (i.e., $h|_{[f]}=const$) for all $f\in\mathcal{F}$.
@@
    Let $Y$ be a manifold, $X=\bigsqcup_{y\in Y}X_{y}=\bigsqcup_{y\in Y}\pi^{-1}(y)$ a fiber bundle over $Y$ with projection $\pi:X\rightarrow Y$. Let $\mathcal{X}:=\{X_{y}\}_{y\in Y}=\{\pi^{-1}(y)\}_{y\in Y}$. Then, in the notation of Definition 3, we get a map on global sections of $X$ given by
@@
    In this case, the limit topology of Definition 3 on $X$ plays a nontrivial role as a topology (induced by that of the fibers $X_{y}$ of the bundle $X$) that can be compared with the underlying topology of $X$.
@@ 4. Concrete quotient-realization and metrization of compact-subset hyperspaces
Definition 12 (Chained space, Topologies of uniform/pointwise/compact-uniform convergence).
A chained-space $X=(X,\mathcal{U})$ is a space $X$ together with a chain $\mathcal{U}=(\{\mathcal{U}_{\lambda}\}_{\lambda\in\Lambda},\preceq)$ of open covers of $X$, where $\mathcal{U}_{\lambda}\preceq\mathcal{U}_{\lambda^{\prime}}$ iff for any $O\in\mathcal{U}_{\lambda}$ there is $O^{\prime}\in\mathcal{U}_{\lambda^{\prime}}$ such that $O\subset O^{\prime}$ and [for any $O^{\prime}\in\mathcal{U}_{\lambda^{\prime}}$ and any $\in\mathcal{U}_{\lambda}, OO^{\prime}$ is not a proper subset of OO]”. A collection of open sets $\mathcal{O}$ of $X$ is homogeneous (with respect to $\mathcal{U}$) if $\mathcal{O}\subset\mathcal{U}_{\lambda}$ for some $\lambda$.
Let $X=(X,\mathcal{U})$ be a chained-space, $Y$ a set, and $\mathcal{F}\subset X^{Y}$. A net $\{f_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{F}$ converges uniformly on $Z\subset Y$ to $f\in\mathcal{F}$, written $f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f$, if for any homogeneous system $\mathcal{O}:=\{O_{z}\ni f(z):O_{z}\in\mathcal{U}_{\lambda},z\in Z\}$ of open sets of $X$ ($\big$(e.g., $\mathcal{O}=\{O\}$ for an open neighborhood $\mathcal{U}_{\lambda}\ni O\supset\overline{f(Z)}$ of $\overline{f(Z)}$ in $X$, or equivalently, $q^{-1}(O^{+})\ni f$ of $f$ in $\mathcal{F}\big)$, there exists $\alpha^{\mathcal{O}}\in\mathcal{A}$ such that for each $z\in Z$, $\{f_{\alpha}(z)\}_{\alpha\geq\alpha^{\mathcal{O}}}\subset O_{z}$ ($\big$(resp., e.g., $\exists\alpha^{O}\in\mathcal{A}$ such that $\bigcup_{\alpha\geq\alpha^{O}}f_{\alpha}(Z)\subset O\big)$. (footnote [^7] We can refer to uniform convergence defined without reference to a chain $\mathcal{U}$ as unconditional-uniform convergence. That is, a net $\{f_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{F}$ converges unconditionally-uniformly on $Z\subset Y$ to $f\in\mathcal{F}$, written $f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f$, if for any system $\mathcal{O}:=\{O_{z}\ni f(z):z\in Z\}$ of open sets in $X$, there exists $\alpha^{\mathcal{O}}\in\mathcal{A}$ such that for each $z\in Z$, $\{f_{\alpha}(z)\}_{\alpha\geq\alpha^{\mathcal{O}}}\subset O_{z}$.
NB: All subsequent concepts/results based on (conditional-) uniform convergence, that do not refer to (hence do not depend on) any details about the chain $\mathcal{U}$, also apply-to/hold-for unconditional-uniform convergence. In particular, Notes 2-4 below (and more) do not depend on $\mathcal{U}$.
@@
If $Z\subset Z^{\prime}\subset Y$, then $f_{\alpha}\stackrel{{\scriptstyle u|_{Z^{\prime}}}}{{\longrightarrow}}f$ implies $f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f$ (which holds because each system of open sets $\mathcal{O}:=\{O_{z}\ni f(z):O_{z}\in\mathcal{U}_{\lambda},z\in Z\}$ can be extended to a system of open sets $\mathcal{O}^{\prime}:=\{O_{z}\ni f(z):O_{z}\in\mathcal{U}_{\lambda},z\in Z^{\prime}\}$). In particular, for any $Z\subset Y$,
$\textrm{$f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f~$ ~implies~ $~f_{\alpha}|_{Z}\stackrel{{\scriptstyle\tau_{p}}}{{\longrightarrow}}f|_{Z}~$ (i.e., $~f_{\alpha}(z)\rightarrow f(z)~$ for all $z\in Z$)}.$
@@
A net $\{f_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{F}$ converges uniformly with respect to a family of sets $\mathcal{Z}\subset\mathcal{P}^{\ast}(Y)$ to $f\in\mathcal{F}$ (making it a $\mathcal{Z}$-uniformly convergent net), written $f_{\alpha}\stackrel{{\scriptstyle u|_{\mathcal{Z}}}}{{\longrightarrow}}f$, if $f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f$ for all $Z\in\mathcal{Z}$. We note that $\{f_{\alpha}\}$ is uniformly convergent if and only if $\{Y\}$-uniformly convergent, if and only if (by Note 2 above) uniformly convergent on $Z$ for every $Z\subset Y$.
@@
A subset $\mathcal{C}\subset X^{Y}$ is $\mathcal{Z}$-uniformly closed if every $\mathcal{Z}$-uniformly convergent net in $\mathcal{C}$ converges to a point in $\mathcal{C}$. The topology of $\mathcal{Z}$-uniform convergence $\tau_{uc|_{\mathcal{Z}}}$ on $X^{Y}$ is the topology whose closed sets are the $\mathcal{Z}$-uniformly closed subsets of $X^{Y}$. When $\mathcal{Z}=\{Y\}$, we simply drop $\mathcal{Z}$ from the terminology, i.e., $\mathcal{Z}$-uniformly closed”, $\tau_{uc|_{\mathcal{Z}}}$ (“topology of $\mathcal{Z}$-uniform convergence”), etc become “uniformly closed”, $\tau_{uc}$ (“topology of uniform convergence”), etc. If $\mathcal{Z}=\big\{\{y\}:y\in Y\big\}$, then $\tau_{uc|_{\mathcal{Z}}}=\tau_{p}$, in which case we replace “uniform” or “uniformly” with “pointwise”. When $Y$ is a space and $\mathcal{Z}=K(Y)$, we call $\tau_{\textrm{cc}}\equiv\tau_{\textrm{cuc}}:=\tau_{uc|_{K(Y)}}$ the topology of compact-uniform convergence (which is a generalization of the usual notion of “compact convergence” or “uniform convergence on compact sets” for a metric $X$).
Note 3.
@@
If $\mathcal{Z}\subset\mathcal{Z}^{\prime}$, then $\tau_{uc|_{\mathcal{Z}}}\subset\tau_{uc|_{\mathcal{Z}^{\prime}}}$. To see this, take a $\tau_{uc|_{\mathcal{Z}}}$-closed set $\mathcal{C}\subset X^{Y}$ and show that it is also $\tau_{uc|_{\mathcal{Z}^{\prime}}}$-closed: Indeed, if $f_{\alpha}\in\mathcal{C}$, then $f_{\alpha}\stackrel{{\scriptstyle\tau_{uc|_{\mathcal{Z}^{\prime}}}}}{{\longrightarrow}}f$ implies $f_{\alpha}\stackrel{{\scriptstyle\tau_{uc|_{\mathcal{Z}}}}}{{\longrightarrow}}f$, and so $f\in\mathcal{C}$.
Note 4.
@@
For any $\mathcal{Z}\subset\mathcal{P}(Y)$, with $Y_{\mathcal{Z}}:=\bigcup\{\mathcal{P}(Z):Z\in\mathcal{Z}\}$, we have $\tau_{uc|_{\mathcal{Z}}}=\tau_{uc|_{\{Y_{\mathcal{Z}}\}}}$. This follows from Note 2 above.
Note 5.
@@
Let $X$ be a metric space, $Y$ a space, and $\mathcal{F}\subset X^{Y}$. In $\mathcal{F}$, if $f_{\alpha}\stackrel{{\scriptstyle\tau_{\textrm{cc}}}}{{\longrightarrow}}f$, then $f_{\alpha}\stackrel{{\scriptstyle d^{K}}}{{\longrightarrow}}f$ (i.e., $\forall\varepsilon>0$, $\exists\alpha^{\varepsilon}$ s.t. $d^{K}(f_{\alpha},f)<\varepsilon$ $\forall\alpha\geq\alpha^{\varepsilon}$) for all $K\in K(Y)$, where $d^{K}(f_{\alpha},f):=\sup_{y\in K}d(f_{\alpha}(y),f(y))$. In particular, in the continuous maps $\mathcal{F}:=C(Y,X)\subset X^{Y}$, $f_{\alpha}\stackrel{{\scriptstyle\tau_{\textrm{cc}}}}{{\longrightarrow}}f$ if and only if $f_{\alpha}\stackrel{{\scriptstyle d^{K}}}{{\longrightarrow}}f$ for every $K\in K(Y)$. (footnote [^8] Proof: ($\Rightarrow$): Assume $f_{\alpha}\stackrel{{\scriptstyle\tau_{\textrm{cc}}}}{{\longrightarrow}}f$. Fix any $K\in K(Y)$. Then for any collection $\mathcal{O}=\{O_{y}\ni f(y):y\in K\}$ of open sets in $X$, there is $\alpha^{\mathcal{O}}$ such that $\{f_{\alpha}(y)\}_{\alpha\geq\alpha^{\mathcal{O}}}\subset O_{y}$ (for all $y\in K$). In particular, for any $\varepsilon>0$, with $\mathcal{O}_{\varepsilon}:=\{O_{y,\varepsilon}:=B_{\varepsilon}^{K}(f(y))~|~y\in K\}$, $\{f_{\alpha}(y)\}_{\alpha\geq\alpha^{\mathcal{O}_{\varepsilon}}}\subset O_{y,\varepsilon}~\forall y\in K$ implies $d^{K}(f_{\alpha},f)=\sup_{y\in K}d(f_{\alpha}(y),f(y))<\varepsilon~~~~\forall\alpha\geq\alpha^{\varepsilon}:=\alpha^{\mathcal{O}_{\varepsilon}}$. ($\Leftarrow$): Assume $f_{\alpha}\stackrel{{\scriptstyle d^{K}}}{{\longrightarrow}}f$ in $C(Y,X)$ for all $K\in K(Y)$. Fix $K\in K(Y)$. For any $\varepsilon>0$, there is $\alpha^{\varepsilon}$ such that $d^{K}(f_{\alpha},f)<\varepsilon$ for all $\alpha\geq\alpha^{\varepsilon}$. Therefore, $\{f_{\alpha}(y)\}_{\alpha\geq\alpha^{\varepsilon}}\subset B_{\varepsilon}(f(y))$ for all $\in K$. Consider any collection $\mathcal{O}=\{O_{y}\ni f(y):y\in K\}$ of open sets in $X$. Let $B_{r_{y}}(f(y))\subset O_{y}$ for each $y\in K$. Then by the compactness of $f(K)$, we can choose $r>0$ (independent of $y$) such that $B_{r}(f(y))\subset O_{y}$ for each $y\in K$. Hence, with $\varepsilon<r$ and $\alpha^{\mathcal{O}}:=\alpha^{\varepsilon}$, we get $\{f_{\alpha}(y)\}_{\alpha\geq\alpha^{\mathcal{O}}}=\{f_{\alpha}(y)\}_{\alpha\geq\alpha^{\varepsilon}}\subset B_{\varepsilon}(f(y))\subset B_{r}(f(y))\subset O_{y}$ for all $y\in K$. ).
@@
$\mathcal{S}\mathcal{B}_{luc}:=\left\{B^{d^{K}}_{r}(f):=\{g\in\mathcal{F}:d^{K}(f,g)<r\}~|~r>0,f\in\mathcal{F},K\in K(Y)\right\}.$
Note 6.
By Note 5 above: (i) On any $\mathcal{F}\subset X^{Y}$,   $\tau_{\textrm{cc}}\supset\tau_{luc}:=\bigcap\{\tau_{d^{K}}:K\in K(Y)\}$,   where a set $O\subset\mathcal{F}$ is $\tau_{luc}$-open (resp., $C\subset\mathcal{F}$ is $\tau_{luc}$-closed) $\iff$ $d^{K}$-open (resp., $d^{K}$-closed) for all $K\in K(Y)$. (ii) On continuous maps $\mathcal{F}:=C(Y,X)\subset X^{Y}$, $\tau_{\textrm{cc}}=\tau_{luc}$.
Definition 13 (An interpretation of the Vietoris topology).
Let $X$ be a space and $A_{\alpha},A\in\mathcal{P}^{\ast}(X)$. We say $A_{\alpha}$ centrally-converges (resp., marginally-converges) to $A$ in $X$, written $A_{\alpha}\stackrel{{\scriptstyle+}}{{\longrightarrow}}A$ (resp., $A_{\alpha}\stackrel{{\scriptstyle-}}{{\longrightarrow}}A$), if for any open set $O\subset X$ containing $A$ (resp., meeting $A$), there exists $\alpha_{+}^{O}$ (resp., $\alpha_{-}^{O}$) such that
@@
that is, if $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}^{+}}}{{\longrightarrow}}A$ (resp., $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}^{-}}}{{\longrightarrow}}A$). A set $\mathcal{C}\subset P^{\ast}(X)$ is centrally-closed (resp., marginally-closed) if for any net $A_{\alpha}\in\mathcal{C}$ such that $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}^{+}}}{{\longrightarrow}}A$ (resp., $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}^{-}}}{{\longrightarrow}}A$), we have $A\in\mathcal{C}$. The upper Vietoris topology $\tau_{v}^{+}$ (resp., lower Vietoris topology $\tau_{v}^{-}$) of $\mathcal{P}^{\ast}(X)$ is the topology whose closed sets are the centrally-closed (resp., marginally-closed) subsets of $\mathcal{P}^{\ast}(X)$. Noting that $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}}}{{\longrightarrow}}A$ if and only if $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}^{+}}}{{\longrightarrow}}A$ and $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}^{-}}}{{\longrightarrow}}A$, the Vietoris topology $\tau_{v}$ of $\mathcal{P}^{\ast}(X)$ is the topology whose closed sets are those subsets of $\mathcal{P}^{\ast}(X)$ that are each both centrally-closed and marginally-closed.